Express the equation: y-4 = -4/5(x + 1) in Slope Intercept Form
Answer by lwsshak3(11628)

Express the equation: y-4 = -4/5(x + 1) in Slope Intercept Form
y-4 = -4/5(x + 1)
lcd:5
5y-20=-4x-4
5y=-4x+16
y=-4x/5+16/5